OpenStudy (anonymous):

please help!! i will medal and fan!! Simplify 7to the power of6 ⋅ 7to the power of5. 49to the power of11 49to the power of35 7to the power of11 7to the power of35

OpenStudy (anonymous):

@aum maybe can you help me again?

OpenStudy (anonymous):

d.

OpenStudy (anonymous):

i think

OpenStudy (aum):

\[\large a^m * a^n = a^{m+n} \\ \large \text{ } \\ 7^6 * 7^5 = 7^{6+5} = ? \]

OpenStudy (anonymous):

what? this is really confusing, loll @aum